For this problem, let f(x) = 3x^2 - 1 and g(x) = 2x + 5.
Find f(g(-4)) and g(f(-4)).
The books answer is 26;99. I'm not sure how to solve this. Please show me how it's done.

Found 2 solutions by jim_thompson5910, solver91311:
Answer by jim_thompson5910(35256)

f(x) = 3x^2 - 1


f(x) = 3(x)^2 - 1


f(g(x)) = 3(g(x))^2 - 1 ... replace every x with g(x)


f(g(x)) = 3(2x + 5)^2 - 1 ... replace the g(x) on the right side with 2x+5


f(g(-4)) = 3(2(-4) + 5)^2 - 1 ... replace every x with -4. Now evaluate


f(g(-4)) = 3(-8 + 5)^2 - 1


f(g(-4)) = 3(-3)^2 - 1


f(g(-4)) = 3(9) - 1


f(g(-4)) = 27 - 1


f(g(-4)) = 26


I'll let you do the other part.

Answer by solver91311(24713)





To find , just substitute -4 in place of and do the arithmetic.

The other problem works the same way.

John

My calculator said it, I believe it, that settles it